OpenStudy (anonymous):

Consider the line that passes through the points (2, 3) and (6, 11). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y -11 = 2(x - 6) y - 3 = 2(x - 2) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two given points.

OpenStudy (anonymous):

simply put the points into the problem if equation 1 # it is a point on that line

OpenStudy (anonymous):

equation #1 when puting in the both points (x,y), you get that those points are on the line

OpenStudy (amistre64):



OpenStudy (amistre64):

as long as your slope is 2, they both look good

OpenStudy (anonymous):

substituting into equation 2 you get the same point so yes both lines pass through the points

OpenStudy (amistre64):

(6, 11) -(2, 3) ------- 4,8 = 8/4 = 2

OpenStudy (amistre64):

i concur :) both are good

OpenStudy (anonymous):

or you can solve for y and see that both are the same exact line y=2x-1 you can just assume that since their slopes are the same that they will hit the same points thats like saying that will hit the same point

OpenStudy (anonymous):

You CANT

OpenStudy (anonymous):

so just finding the slope would not help you'd have to make sure the lines are the same other wise the slope being the same just means that they are parallel and will never meet the same points